If sin A + sin B = 2
Find cos2 A+ cos2B
Find the  value of the container 
Thank you

Grade:10

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

sin A + sin B = 2

implies A, B = (pi/2)

therefore (cos A)^2 + ​(cos B)^2 = 0+0 = 0